A Quiz!
A Farmer Plants a Tree in a Square Plot of Land and
finds that it is 100 m and 140 m from two opposite
corners and 60 m from the 3rd corner.
What is the area of the plot?

Answer by Alan3354(69443)

A Farmer Plants a Tree in a Square Plot of Land and
finds that it is 100 m and 140 m from two opposite
corners and 60 m from the 3rd corner.
What is the area of the plot?
--------------------------------
Lines from the tree to the corners of the square make 3 scalene triangles.
s = the sides of the square.
A, B & C are the angles at the tree.
Use the Cosine Law.
----
s^2 = 60^2 + 140^2 - 2*60*140*cos(A)
s^2 = 60^2 + 100^2 - 2*60*100*cos(B)
2s^2 = 100^2 + 140^2 - 2*100*140*cos(C)
A + B + C = 360
----
s^2 = 23200 - 16800*cos(A) = 13600 - 12000*cos(B) = 14800 - 14000*cos(C)
-----------------------
Can you do the rest?